Which equation is an identity?
a. 11-(2v+3)=-2v-8
b. 5w+8-w=6w-2(w-4)
c. 7m-2=8m+4-m
d. 8y+9=8y-3

Answer by Edwin McCravy(20054)

a. 11 - (2v + 3) = -2v - 8

Simplify

     11 - 2v - 6 = -2v - 8

          5 - 2v = -2v - 8

Add 3v to both sides

               5 = -8

That's false, so it is not an identity.




b. 5w + 8 - w = 6w - 2(w - 4) 
Simplify

       4w + 8 = 6w - 2w + 8

       4w + 8 = 4w + 8

That's exactly the same on both sides, so it's
an identity.

        
c. 7m - 2 = 8m + 4 - m

   7m - 2 = 7m + 4

Subtract 7m from both sides:

       -2 = 4

That's false, so it is not an identity.
  

d. 8y + 9 = 8y - 3

Subtract 8y from both sides:

        9 = -3

That's false, so it is not an identity.

Edwin